OpenStudy (anonymous):

a leaky faucet fills a 100 milileter cup in 10 min. how much water is wasted in a day, week and year???? how long does it take to wast a kiloliter of water????

OpenStudy (anonymous):

100 ml in 10 min = 10 ml/m * 60 m / h * 24 h/d = 14,400 ml in a day 14400 * 7 d/w = 100,800 ml in a week 100,800 * 52 w / y = 5,241,600 ml in a year to waste one kilometer would be to waste ONE MILLION ML. 100,800 in one week, so if X = # of weeks, we have 100800*X = 1,000,000 x = 9.92 weeks

OpenStudy (anonymous):

wow thats a lot of numbers

OpenStudy (anonymous):

you asked for it :)

OpenStudy (anonymous):

is there any possible way to dumb it down a lottle bit more

OpenStudy (anonymous):

m = minutes h = hours d = days w = weeks y = years seems pretty standard. those are the answers, no way around it.

OpenStudy (anonymous):

sorry im in the lower math class so thanks

OpenStudy (anonymous):

it isn't an easy problem. make sure you understand it before you hand it in or UH OHH

OpenStudy (anonymous):

yea my teacher is kinda clueless so she wont really be aware that like no one gets it

OpenStudy (anonymous):

ok lol why don't u ask her to explain it. thats what a teacher is for.

OpenStudy (anonymous):

yea like i said shes clueless it doesnt help

OpenStudy (anonymous):

bu thanks again

OpenStudy (anonymous):

anytime man